# Loss of integrability in a system with two-photon interactions

### v1

@@ I Introduction
Quantum systems involving the interaction between light and matter play a crucial role in a wide range of quantum technologies Dowling and Milburn (2003), including quantum information and communication Zeilinger (1999); Gisin and Thew (2007); Swingle et al. (2016); Landsman et al. (2019), quantum metrology and sensing Taylor and Bowen (2016); Degen et al. (2017); Fiderer and Braun (2018); Crawford et al. (2021); Montenegro et al. (2024), and quantum simulation Lloyd (1996); Georgescu et al. (2014), among others. The fundamental theoretical framework for describing such interactions is the Rabi model Rabi (1936, 1937); Braak et al. (2016), which considers a single two-level atom or qubit coupled to a single mode of a quantized electromagnetic field, allowing for the creation or annihilation of a single photon at a time. Extensions of the Rabi model to multi-qubit systems have led to the Dicke model Dicke (1954); Kirton et al. (2019); Roses and Dalla Torre (2020); Villaseñor et al. (2024).
Integrable versions of the Rabi and Dicke models arise when counter-rotating terms in the Hamiltonian are neglected under the rotating wave approximation (RWA) Scully and Zubairy (1997). This simplification yields the Jaynes-Cummings Jaynes and Cummings (1963); Larson and Mavrogordatos (2021) and Tavis–Cummings models Tavis and Cummings (1968); Garraway (2011), which correspond to the integrable single-qubit and multi-qubit cases, respectively.
@@
Subsequent investigations have further examined solvability aspects of the two-photon Rabi model and its variants Peng et al. (2012, 2013); Maciejewski and Stachowiak (2017); Cui et al. (2017), as well as quantum dynamical properties Lü et al. (2017). Although generalizations to multi-photon interactions have been proposed, it has been argued that processes involving more than two photons are unphysical in the Rabi model Lo et al. (1998); Braak (2025). Nevertheless, recent findings highlight the potential of the two-photon Rabi model for quantum metrology applications Ying et al. (2025); Ying (2025) and outline efficient schemes for generating two-photon processes in the presence of dissipation Mazhari and Rituraj (2025).
@@
When extended to multi-qubit systems, the two-photon Rabi model leads to the two-photon Dicke model, which also exhibits spectral collapse Felicetti et al. (2015); Lo (2021). Research has investigated phenomena such as photon squeezing in this model Banerjee et al. (2022) and its integrable limit Gerry and Togeas (1989), known as the two-photon Tavis–Cummings model. In the single-photon Dicke model, both ground-state and excited-state quantum phase transitions are well known Emary and Brandes (2003a, b); Lambert et al. (2004); Brandes (2005, 2013), giving rise to a superradiant phase marked by a significant increase in the average photon number. A similar collective superradiant phase has been identified in the two-photon Dicke model in the ultrastrong coupling regime Garbe et al. (2017).
@@
Interestingly, multi-photon interactions beyond two photons are possible in the two-photon Dicke model due to the presence of special solutions known as dark states Peng et al. (2017), distinguishing it from the two-photon Rabi model, where such interactions are forbidden Lo et al. (1998); Braak (2025). Further studies have also incorporated Stark shift terms into the model Klimov et al. (1999); Zhai et al. (2025). Experimental realizations of two-photon processes date back to early investigations of two-photon lasing Loy (1978); Schlemmer et al. (1980); Nikolaus et al. (1981); Gauthier et al. (1992) and masers Brune et al. (1987a, b, 1988), and have since been explored in systems such as quantum dots Stufler et al. (2006); del Valle et al. (2010) and Rydberg atoms Bertet et al. (2002); Zhang et al. (2013). More recently, two-photon processes have been proposed for the development of quantum batteries Crescente et al. (2020).
@@
Dissipative effects have also been studied within the two-photon Dicke model, particularly in relation to phase transitions Garbe et al. (2020); Shah et al. (2024) and chaotic dynamics Li et al. (2022); Li and Chesi (2024). While progress has been made in understanding dissipative chaos, less attention has been given to the emergence of chaos in the isolated (non-dissipative) two-photon Dicke model Wang et al. (2019). In this work, we investigate the loss of integrability in the isolated two-photon Dicke model and characterize the onset of chaos in its normal phase using both classical and quantum diagnostics. Our results provide new insights into the nature of integrable, mixed, and fully chaotic regimes in this system.
@@
The remainder of this article is organized as follows. In Sec. II, we introduce the two-photon Dicke model, discuss its general properties, and derive its classical limit in the normal phase via a mean-field approximation. Section III is devoted to analyzing the integrable regime of the model, where we obtain analytical expressions for the eigenvalues and eigenstates using a Bogoliubov transformation, which are employed to characterize the loss of integrability in the low-energy regime. In Sec. IV, we investigate the breakdown of integrability and characterize the onset of chaos in the normal phase through both classical and quantum indicators. Finally, in Sec. V, we summarize our main results and present our conclusions. Additional technical details and derivations are provided in the appendices.
@@ II Two-Photon Dicke model
$\hat{H}_{\text{D}}=\omega\hat{a}^{\dagger}\hat{a}+\omega_{0}\hat{J}_{z}+\frac{\gamma}{\mathcal{N}}\left(\hat{a}^{\dagger 2}+\hat{a}^{2}\right)\left(\hat{J}_{+}+\hat{J}_{-}\right),$ (1)
where the operators $\hat{a}^{\dagger}$ and $\hat{a}$ are the creation and annihilation operators of the bosonic mode which satisfy the commutation relation $[\hat{a},\hat{a}^{\dagger}]=\hat{\mathbb{I}}$. The operators $\hat{J}_{z}$ and $\hat{J}_{\pm}=\hat{J}_{x}\pm i\hat{J}_{y}$ are collective pseudospin operators, $\hat{J}_{x,y,z}=\frac{1}{2}\sum_{k=1}^{\mathcal{N}}\hat{\sigma}_{x,y,z}^{k}$, that obey the SU(2) algebra of the Pauli matrices $\hat{\sigma}_{x,y,z}$. The parameters $\omega$ and $\omega_{0}$ represent the field frequency and the atomic level splitting, respectively. The parameter $\gamma$ is the atom-photon coupling strength, whose critical value $\gamma_{\text{sc}}=\omega/2$ identifies the spectral collapse Felicetti et al. (2015); Duan et al. (2016); Rico et al. (2020).
The two-photon Dicke Hamiltonian commutes with the parity operator $\hat{\Pi}=e^{i\pi\hat{\Lambda}}$, $[\hat{H}_{\text{D}},\hat{\Pi}]=0$, where the operator $\hat{\Lambda}=\hat{a}^{\dagger}\hat{a}/2+\hat{J}_{z}+j\hat{\mathbb{I}}$ defines the generalized number of excitations. This parity symmetry corresponds to the simultaneous exchange of the operators $\hat{a}\to i\hat{a},\hat{a}^{\dagger}\to-i\hat{a}^{\dagger}$ and $\hat{J}_{x}\to-\hat{J}_{x}$, and divides the system into four subspaces identified by the eigenvalues of the parity operator $p=\pm 1,\pm i$. The Hamiltonian also commutes with the squared total pseudospin operator, $[\hat{H}_{\text{D}},\hat{\mathbf{J}}^{2}]=0$, whose eigenvalues $j(j+1)$ specify different subspaces. In this work, we study the case with $j=\mathcal{N}/2$.
The ground state of the two-photon Dicke model undergoes a phase transition from a normal phase, characterized by the absence of photons and atomic excitations, to a superradiant phase, in which both quantities attain finite values Garbe et al. (2017); Banerjee et al. (2022). The majority of the results presented in this work pertain to the normal phase.
@@ II.1 Classical two-photon Dicke Hamiltonian in the normal phase
We employ a mean-field approximation using coherent states to derive the classical limit of the two-photon Dicke model in the normal phase. This approach follows the standard procedure established for the single-photon Dicke model de Aguiar et al. (1992); Bakemeier et al. (2013); Bastarrachea-Magnani et al. (2014a, b, 2015); Chávez-Carlos et al. (2016). The complete derivation is presented in Appendix A. We use Glauber coherent states
@@
associated to the bosonic subspace and Bloch coherent states
@@
associated to the atomic subspace, to define an overall Glauber-Bloch coherent state as the tensor product
@@
where $\mathbf{x}=(q,p;Q,P)$ are canonical position and momentum classical variables associated with the bosonic and atomic phase space through the parameters $\alpha(q,p)$ and $\beta(Q,P)$, respectively. The last parameters are defined in Eq. (18) and Eq. (19).
@@
appropriately scaled by the system size $j$. This scaling yields an intensive classical energy shell $\epsilon=E/j$ and defines an effective Planck constant $\hbar_{\text{eff}}=1/j$ Ribeiro et al. (2006), which facilitates the semiclassical analysis.
@@ III Integrability of the two-photon Dicke Hamiltonian
Figure 1: Peres lattices of the operators [(a1)-(c1)] $\hat{c}^{\dagger}\hat{c}$, [(a2)-(c2)] $\hat{J}_{x}$, and [(a3)-(c3)] $\hat{J}_{x}^{2}$. Each column identifies a different coupling strength: [(a1)-(a3)] $\gamma=0.2$, [(b1)-(b3)] $\gamma=0.3$, and [(c1)-(c3)] $\gamma=0.4$. In each panel, the black dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian in Eq. (1) and the red lines describe the analytical values using Eq. (10). In panels (a1)-(c1) and (a2)-(c2), each red line represents $m_{x}$ as a continuous variable for discrete values of $n_{c}$. In panels (a3)-(c3), the continuous variable is $m_{x}^{2}$. System parameters: $\omega=1$, $\omega_{0}=0$, and $j=15$. We use a perturbation parameter $\varepsilon=0.001$. The truncation value is $n_{\text{max}}=200$.
@@ III.1 Integrable limit of zero atomic level splitting
where $\Omega_{m_{x}}\equiv\omega\sqrt{1-4\lambda_{m_{x}}^{2}}$ and $\lambda_{m_{x}}\equiv 2\gamma m_{x}/(\omega\mathcal{N})$. The operators $\hat{c}$ and $\hat{c}^{\dagger}$ satisfy the commutation relation $[\hat{c},\hat{c}^{{\dagger}}]=\hat{\mathbb{I}}$ and are defined through the modified annihilation operator
@@ III.2 Integrable correspondence
A useful method for visually identifying the loss of integrability in a quantum system is the analysis of the energy-dependent expectation value of a chosen operator, $\langle E_{k}|\hat{O}|E_{k}\rangle$, a technique known as the Peres lattice Peres (1984); Feingold et al. (1985); Feingold and Peres (1986). In this work, we compute the Peres lattices for the operators $\hat{c}^{\dagger}\hat{c}$, $\hat{J}_{x}$, and $\hat{J}_{x}^{2}$. As the pseudospin operator $\hat{J}_{x}$ does not commute with the system’s parity operator, i.e., $[\hat{J}_{x},\hat{\Pi}]\neq 0$, its expectation value vanishes for eigenstates of definite parity due to orthogonality between states of different parity. To circumvent this and compute meaningful expectation values of $\hat{J}_{x}$, we explicitly break the parity symmetry by adding a small perturbative term to the Hamiltonian in Eq. (20), namely $\hat{H}^{\prime}_{\text{D}}+\varepsilon\hat{J}_{x}$, where $\varepsilon$ is a small parameter controlling the strength of the perturbation.
In Fig. 1, we plot the expectation values (black points) of the operators $\hat{c}^{\dagger}\hat{c}$, $\hat{J}_{x}$, and $\hat{J}_{x}^{2}$, which were numerically computed with Eqs. (11)-(13), for a set of increasing couplings $\gamma$. In the same figure, we plot the analytical values (red lines) obtained with Eq. (10). In Figs. 1(a1)-1(c1) and Figs. 1(a2)-1(c2), we display the value $m_{x}$ as a continuous variable (red lines) for discrete values of $n_{c}$. In Figs. 1(a3)-1(c3), the continuous variable is represented by $m_{x}^{2}$. As expected, in all panels, the numerical results are in exact agreement with the analytical predictions.
@@
Figure 2: Peres lattices of the operator $\hat{J}_{x}^{2}$. In each panel, the black dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian in Eq. (1) and the red lines describe the analytical values using Eq. (10). In panel (a), we set $\omega_{0}=0$ with $\varepsilon=0.001$ and plot $m_{x}^{2}$ as a continuous variable for discrete values of $n_{c}$. In panels (b) and (c), we set $\omega_{0}=0.05$ to compute the numerical expectation values (black dots), but we use the analytical results (red lines) from panel (a). In panel (b), the continuous variable is $m_{x}^{2}$ as described in panel (a), while in panel (c), the continuous variable is $n_{c}$ for discrete values of $m_{x}^{2}$. System parameters: $\omega=1$, $\gamma=0.3$, and $j=15$. The truncation value is $n_{\text{max}}=200$.
Figure 3: Peres lattices of the operators (a) $\hat{a}^{\dagger}\hat{a}$, (b) $\hat{J}_{z}$, and (c) $\hat{J}_{x}^{2}$. In each panel, the green and blue dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian in Eq. (1). The blue dots correspond to eigenstates that have a larger projection in the basis of $\hat{J}_{x}$. The green dots correspond to eigenstates that have a larger projection in the basis of $\hat{J}_{z}$. System parameters: $\omega=1$, $\omega_{0}=0.05$, $\gamma=0.3$, and $j=25$. The truncation value is $n_{\text{max}}=200$.

### v2

@@ I Introduction
Quantum systems involving the interaction between light and matter play a crucial role in a wide range of quantum technologies Dowling and Milburn (2003), including quantum information and communication Zeilinger (1999); Gisin and Thew (2007); Swingle et al. (2016); Landsman et al. (2019), quantum metrology and sensing Taylor and Bowen (2016); Degen et al. (2017); Fiderer and Braun (2018); Crawford et al. (2021); Montenegro et al. (2025), and quantum simulation Lloyd (1996); Georgescu et al. (2014), among others. The fundamental theoretical framework for describing such interactions is the Rabi model Rabi (1936, 1937); Braak et al. (2016), which considers a single two-level atom or qubit coupled to a single mode of a quantized electromagnetic field, allowing for the creation or annihilation of a single photon at a time. Extensions of the Rabi model to multiqubit systems have led to the Dicke model Dicke (1954); Kirton et al. (2019); Roses and Dalla Torre (2020); Villaseñor et al. (2024).
Integrable versions of the Rabi and Dicke models arise when counter-rotating terms in the Hamiltonian are neglected under the rotating wave approximation (RWA) Scully and Zubairy (1997). This simplification yields the Jaynes-Cummings Jaynes and Cummings (1963); Larson and Mavrogordatos (2021) and Tavis–Cummings models Tavis and Cummings (1968); Garraway (2011), which correspond to the integrable single-qubit and multiqubit cases, respectively.
@@
Subsequent investigations have further examined solvability aspects of the two-photon Rabi model and its variants Peng et al. (2012, 2013); Maciejewski and Stachowiak (2017); Cui et al. (2017), as well as quantum dynamical properties et al. (2017). Although generalizations to multiphoton interactions have been proposed, it has been argued that processes involving more than two photons are unphysical in the Rabi model Lo et al. (1998); Braak (2026). Nevertheless, recent findings highlight the potential of the two-photon Rabi model for quantum metrology applications Ying et al. (2025); Ying (2025) and outline efficient schemes for generating two-photon processes in the presence of dissipation Mazhari and Rituraj (2025).
@@
When extended to multiqubit systems, the two-photon Rabi model leads to the two-photon Dicke model, which also exhibits spectral collapse Felicetti et al. (2015); Lo (2021). Research has investigated phenomena such as photon squeezing in this model Banerjee et al. (2022) and its integrable limit Gerry and Togeas (1989), known as the two-photon Tavis–Cummings model. In the single-photon Dicke model, both ground-state and excited-state quantum phase transitions are well known Emary and Brandes (2003a, b); Lambert et al. (2004); Brandes (2005, 2013), giving rise to a superradiant phase marked by a significant increase in the average photon number. A similar collective superradiant phase has been identified in the two-photon Dicke model in the ultrastrong coupling regime Garbe et al. (2017).
@@
Interestingly, multiphoton interactions beyond two photons are possible in the two-photon Dicke model due to the presence of special solutions known as dark states Peng et al. (2017), distinguishing it from the two-photon Rabi model, where such interactions are forbidden Lo et al. (1998); Braak (2026). Further studies have also incorporated Stark shift terms into the model Klimov et al. (1999); Zhai et al. (2025). Experimental realizations of two-photon processes date back to early investigations of two-photon lasing Loy (1978); Schlemmer et al. (1980); Nikolaus et al. (1981); Gauthier et al. (1992) and masers Brune et al. (1987a, b, 1988), and have since been explored in systems such as quantum dots Stufler et al. (2006); del Valle et al. (2010) and Rydberg atoms Bertet et al. (2002); Zhang et al. (2013). More recently, two-photon processes have been proposed for the development of quantum batteries Crescente et al. (2020).
@@
Dissipative effects have also been studied within the two-photon Dicke model, particularly in relation to phase transitions Garbe et al. (2020); Jayesh Shah et al. (2025) and chaotic dynamics Li et al. (2022); Li and Chesi (2024). While progress has been made in understanding dissipative chaos, less attention has been given to the emergence of chaos in the isolated (nondissipative) two-photon Dicke model Wang et al. (2019). In this work, we investigate the loss of integrability in the isolated two-photon Dicke model and characterize the onset of chaos in its normal phase using both classical and quantum diagnostics. Our results provide new insights into the nature of integrable, mixed, and fully chaotic regimes in this system.
@@
The remainder of this article is organized as follows: In Sec. II, we introduce the two-photon Dicke model, discuss its general properties, and derive its classical limit in the normal phase via a mean-field approximation with coherent states. Section III is devoted to analyzing the integrable regime of the model, where we obtain analytical expressions for the eigenvalues and eigenstates using a Bogoliubov transformation, which are employed to characterize the loss of integrability in the low-energy regime. In Sec. IV, we investigate the breakdown of integrability and characterize the onset of chaos in the normal phase through both classical and quantum indicators. Finally, in Sec. V, we summarize our main results and present our conclusions. Additional technical details and derivations are provided in the appendices.
@@ II Two-Photon Dicke model
[…]
where the operators $\hat{a}^{\dagger}$ and $\hat{a}$ are the creation and annihilation operators of the bosonic mode which satisfy the commutation relation $\hat{a},\hat{a}^{\dagger}]=\hat{\mathbb{I}}$. The operators $\hat{J}_{z}$ and $\hat{J}_{\pm}=\hat{J}_{x}\pm i\hat{J}_{y}$ are collective pseudospin operators, $\hat{J}_{x,y,z}=\frac{1}{2}\sum_{k=1}^{\mathcal{N}}\hat{\sigma}_{x,y,z}^{k}$, that obey the SU(2) algebra of the Pauli matrices $\hat{\sigma}_{x,y,z}$. The parameters $\omega$ and $\omega_{0}$ represent the field frequency and the atomic level splitting, respectively. The parameter $\gamma$ is the atom-photon coupling strength, whose critical value $\gamma_{\text{sc}}=\omega/2$ identifies the spectral collapse Felicetti et al. (2015); Duan et al. (2016); Rico et al. (2020).
The two-photon Dicke Hamiltonian commutes with the parity operator $\hat{\Pi}=e^{i\pi\hat{\Lambda}}$, $\hat{H}_{\text{D}},\hat{\Pi}]=0$, where the operator $\hat{\Lambda}=\hat{a}^{\dagger}\hat{a}/2+\hat{J}_{z}+j\hat{\mathbb{I}}$ defines the generalized number of excitations. This parity symmetry corresponds to the simultaneous exchange of the operators $\hat{a}\to i\hat{a},\hat{a}^{\dagger}\to-i\hat{a}^{\dagger}$, and $\hat{J}_{x}\to-\hat{J}_{x}$ and divides the system into four subspaces identified by the eigenvalues of the parity operator $p=\pm 1,\pm i$. The Hamiltonian also commutes with the squared total pseudospin operator, $\hat{H}_{\text{D}},\hat{\mathbf{J}}^{2}]=0$, whose eigenvalues $j(j+1)$ specify different subspaces. In this work, we study the case with $j=\mathcal{N}/2$.
The ground state of the two-photon Dicke model undergoes a phase transition from a normal phase, characterized by the absence of photons and atomic excitations, to a superradiant phase, in which both quantities attain finite values Garbe et al. (2017); Banerjee et al. (2022). The superradiant ground state is well described by a squeezed state. The critical value of the coupling strength that separates both phases is $\gamma_{\text{c}}=\sqrt{\omega\omega_{0}j/2}$  Garbe et al. (2017). The results presented in this work pertain to the normal phase ($\gamma<\gamma_{\text{c}}$) and consider coupling strength values below the spectral collapse threshold ($\gamma<\omega/2$).
@@ II.1 Classical two-photon Dicke Hamiltonian in the normal phase
Previous research has demonstrated that squeezed states effectively describe the ground-state energy of the two-photon Dicke Hamiltonian in the normal and superradiant phase when using a mean-field approximation Garbe et al. (2017). As mentioned above, there is squeezing only in the superradiant phase.
In this work, we apply a mean-field approximation using coherent states to derive the classical limit of the two-photon Dicke model in the normal phase. It accurately provides the ground-state energy and describes the complete classical dynamics in the normal phase of the system without dependency on the system size. The mean-field approximation follows the standard procedure established for the single-photon Dicke model de Aguiar et al. (1992); Bakemeier et al. (2013); Bastarrachea-Magnani et al. (2014a, b, 2015); Chávez-Carlos et al. (2016). The complete derivation is presented in Appendix A. We use Glauber coherent states
@@
associated with the bosonic subspace and Bloch coherent states
@@
associated with the atomic subspace, to define an overall Glauber-Bloch coherent state as the tensor product
@@
where $\mathbf{x}=(q,p;Q,P)$ are canonical position and momentum classical variables associated with the bosonic and atomic phase space through the parameters $\alpha(q,p)$ and $\beta(Q,P)$, respectively. The last parameters are defined in Eqs. (19) and (20).
@@
appropriately scaled by the system size $j$. This scaling yields an intensive classical energy shell
$\epsilon\equiv\frac{E}{j}$ (6)
and defines an effective Planck constant $\hbar_{\text{eff}}=1/j$ Ribeiro et al. (2006), which facilitates the semiclassical analysis.
@@ III Integrability of the two-photon Dicke Hamiltonian
Figure 1: Peres lattices of the operators (a1)-(c1) $\hat{c}^{\dagger}\hat{c}$, (a2)-(c2) $\hat{J}_{x}$, and (a3)-(c3) $\hat{J}_{x}^{2}$, as a function of the scaled energy $\epsilon=E/j$. Each column identifies a different coupling strength: (a1)-(a3) $\gamma=0.2$, (b1)-(b3) $\gamma=0.3$, and (c1)-(c3) $\gamma=0.4$. In each panel, the black dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian in Eq. (1) and the red lines describe the analytical values using Eq. (11). In panels (a1)-(c1) and (a2)-(c2), each red line represents $m_{x}$ as a continuous variable for discrete values of $n_{c}$. In panels (a3)-(c3), the continuous variable is $m_{x}^{2}$. System parameters: $\omega=1$, $\omega_{0}=0$, and $j=15$. We use a perturbation parameter $\varepsilon=0.001$. The truncation value is $n_{\text{max}}=200$.
@@ III.1 Integrable limit of zero atomic level splitting
where $\Omega_{m_{x}}\equiv\omega\left(1-4\lambda_{m_{x}}^{2}\right)^{1/2}$ and $\lambda_{m_{x}}\equiv 2\gamma m_{x}/(\omega\mathcal{N})$. The operators $\hat{c}$ and $\hat{c}^{\dagger}$ satisfy the commutation relation $\hat{c},\hat{c}^{{\dagger}}]=\hat{\mathbb{I}}$ and are defined through the modified annihilation operator
@@ III.2 Integrable correspondence
A useful method for visually identifying the loss of integrability in a quantum system is the analysis of the energy-dependent expectation value of a chosen operator, $\langle E_{k}|\hat{O}|E_{k}\rangle$, a technique known as the Peres lattice Peres (1984); Feingold et al. (1985); Feingold and Peres (1986). In this work, we compute the Peres lattices for the operators $\hat{c}^{\dagger}\hat{c}$, $\hat{J}_{x}$, and $\hat{J}_{x}^{2}$, as a function of the scaled eigenenergies $\epsilon_{k}=E_{k}/j$. As the pseudospin operator $\hat{J}_{x}$ does not commute with the system’s parity operator, i.e., $\hat{J}_{x},\hat{\Pi}]\neq 0$, its expectation value vanishes for eigenstates of definite parity due to orthogonality between states of different parity. To circumvent this and compute meaningful expectation values of $\hat{J}_{x}$, we explicitly break the parity symmetry by adding a small perturbative term to the Hamiltonian in Eq. (22), namely $\hat{H}^{\prime}_{\text{D}}+\varepsilon\hat{J}_{x}$, where $\varepsilon$ is a small parameter controlling the strength of the perturbation.
In Fig. 1, we plot the expectation values (black points) of the operators $\hat{c}^{\dagger}\hat{c}$, $\hat{J}_{x}$, and $\hat{J}_{x}^{2}$, which were numerically computed with Eqs. (12)-(14), for a set of increasing couplings $\gamma$. In the same figure, we plot the analytical values (red lines) obtained with Eq. (11). In Figs. 1(a1)-1(c1) and 1(a2)-1(c2), we display the value $m_{x}$ as a continuous variable (red lines) for discrete values of $n_{c}$. In Figs. 1(a3)-1(c3), the continuous variable is represented by $m_{x}^{2}$. As expected, in all panels, the numerical results are in exact agreement with the analytical predictions.
@@
Figure 2: Peres lattices of the operator $\hat{J}_{x}^{2}$ as a function of the scaled energy $\epsilon=E/j$. In each panel, the black dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian in Eq. (1) and the red lines describe the analytical values using Eq. (11). In panel (a), we set $\omega_{0}=0$ with $\varepsilon=0.001$ and plot $m_{x}^{2}$ as a continuous variable for discrete values of $n_{c}$. In panels (b) and (c), we set $\omega_{0}=0.05$ to compute the numerical expectation values (black dots), but we use the analytical results (red lines) from panel (a). In panel (b), the continuous variable is $m_{x}^{2}$ as described in panel (a), while in panel (c), the continuous variable is $n_{c}$ for discrete values of $m_{x}^{2}$. System parameters: $\omega=1$, $\gamma=0.3$, and $j=15$. The truncation value is $n_{\text{max}}=200$.
Figure 3: Peres lattices of the operators (a) $\hat{a}^{\dagger}\hat{a}$, (b) $\hat{J}_{z}$, and (c) $\hat{J}_{x}^{2}$, as a function of the scaled energy $\epsilon=E/j$. In each panel, the green and blue dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian in Eq. (1). The blue dots correspond to eigenstates that tend to behave as eigentates of the operator $\hat{J}_{x}$. The green dots correspond to eigenstates that tend to behave as eigenstates of the operator $\hat{J}_{z}$. System parameters: $\omega=1$, $\omega_{0}=0.05$, $\gamma=0.3$, and $j=25$. The truncation value is $n_{\text{max}}=200$.